A right triangle has a perimeter of 36 meters, a hypotenuse of 15 meters, and a leg of 9 meters. Find the area of the triangle.
Answer by jorel1380(3719)

If the hypotenuse is 15 meters, and one leg is 9 meters, then the third leg is 36-24 or 12 meters. The area equals 1/2(leg 1 X leg 2)=1/2(12X9)=54 sq. mtrs..